OpenStudy (anonymous):

4x-6y=2 5x+3y=1 Solve the system using any algebraic method

OpenStudy (anonymous):

Use the Elimination Method. If you look at the y's you will see that if you multiply the second equation by the first you will have opposites, so multiply the second equation by 2.

OpenStudy (anonymous):

why multiply by 2?

OpenStudy (anonymous):

\[4x-6y=2\] \[10x+6y=2\]

OpenStudy (anonymous):

to use elimination you need to have the numbers in front of one of the variables either the same or with just different signs, and the best way to do that is to multiply the second equation by 2, make sense?

OpenStudy (anonymous):

And so I get 10x+6y=2? y= 1/7 right?

OpenStudy (anonymous):

oh i see okay :p

OpenStudy (anonymous):

so now you are going to add together the two equations and by doing this the y will go away

OpenStudy (anonymous):

\[14x=4\]

OpenStudy (anonymous):

and you divide both sides by 14, correct?

OpenStudy (anonymous):

yeah, and so x will be 1/7, not y.

OpenStudy (anonymous):

to find y you just plug in x=1/7 to either of the original equations and solve for y

OpenStudy (anonymous):

okay I get it now, thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

I have a similar problem to that, do I do the same thing to it also

OpenStudy (anonymous):

The problem is 4x+6y=8 2x+3y=3

OpenStudy (anonymous):

Ok, so you need to look at the numbers in front of the letters. Is there any pair that you can multiply one of the numbers to get the other number?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

to cancel out the x

OpenStudy (anonymous):

Good, so that would be multiplying the second equestion by -2. Go ahead, this one is going to be tricky, but don't worry.

OpenStudy (anonymous):

I got 4x+6y=8 -4x+3y=-6, so would it be no solution

OpenStudy (anonymous):

You are correct that it would be no solution, you have a small error on your second equation, it should be -6y, incase you need to show your work.

OpenStudy (anonymous):

yea thats what I meant to put, I have to go now, but I would really like it if you could message me or something so you could help me with all the other problems

OpenStudy (anonymous):

You could follow me, i am not on that often now that it is summer.